OpenStudy (anonymous):

Tan ( theta/2 + pi/3) = 1

OpenStudy (anonymous):

I got it guys. Thanks for helping

OpenStudy (jack1):

...sarcasm...? or did u really get it?

OpenStudy (jack1):

if Tan ( theta/2 + pi/3) = 1 ( theta/2 + pi/3) = arctan (1) theta/2 + pi/3 = pi/4 (which is aka 45 degrees) theta/2 + pi/3 - pi/3 = pi/4 - pi/3 theta/2 = 3pi/12 - 4pi/12 theta/2 = - pi/12 so: theta = -2pi/12 = -pi/6 @Navhar

OpenStudy (anonymous):

That is what I got

OpenStudy (anonymous):

Thanks jack1

OpenStudy (jack1):

noice, props nav 's slaters ;D

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."